mathslover (mathslover):

A man fires a bullet of mass 100 g at a speed of 75 ms–1. The gun is of 3 kg mass. What is the recoil velocity of the gun?

mathslover (mathslover):

@Callisto and @hartnn @Xishem

mathslover (mathslover):

oh k so I did like this: given \(m_1\) = 0.1 kg. and \(m_2 \) = 3 kg. u = \(75 ms^{-1}\) recoil velocity = v 0 = 0.1 * 75 +3 v 0 = 7.5 + 3v -3v = 7.5 v = \(\frac{7.5}{3}=\frac{75}{30}=\frac{15}{6}=2.5 ms^{-1}\)

mathslover (mathslover):

hence recoil velocity = 2.5 \(ms^{-1}\)

mathslover (mathslover):

M i right? if yes then why does the book says the answer as "0.25 m/s" ?

OpenStudy (anonymous):

yea~ using the momentum equation m1v1=m2v2 so 0.1(75)=3v2 v2=7.5/3 =2.5 m/s probably the answer is wrong ;)

OpenStudy (ghazi):

yes you're right

OpenStudy (ghazi):

just use conservation of momentum...momentum is not conserved ideally though

mathslover (mathslover):

hmn so the book is wrong?

OpenStudy (ghazi):

for sure..coz i can't see any fault in this calculation...what does your book say?

mathslover (mathslover):

0.25 m/s

OpenStudy (anonymous):

@mathslover your answer is correct.

OpenStudy (ghazi):

nope it will be 2.5

mathslover (mathslover):

thanks a lot all : @ghazi @Kystal @sami-21

OpenStudy (ghazi):

:) YW

OpenStudy (anonymous):

V = -mv/M

OpenStudy (anonymous):

= -0.1*75/3 = -2.5m/s

OpenStudy (anonymous):

-ve sign rep the velocity is the opp direction..))

OpenStudy (xishem):

It sure seems like your book is wrong a lot :P.